OpenStudy (anonymous):

lim as x approaches -2 { f(x)/x^2} find the lim as x approaches -2 f(X)

OpenStudy (turingtest):

\[\lim_{x\to-2}{f(x)\over x^2}\]\[\lim_{x\to-2}f(x)=?\]is this the problem? does it not give a value for the first limit?

OpenStudy (anonymous):

= 1 sorry

OpenStudy (turingtest):

\[\lim_{x\to-2}{f(x)\over x^2}=1\]\[\lim_{x\to-2}f(x)=?\]

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

\[\lim_{x\to-2}{f(x)\over x^2}=\lim_{x\to-2}f(x)\cdot\lim_{x\to-2}\frac1{x^2}=1\]now you can take the limit, which implies that \(\lim_{x\to-2}f(x)=?\)